Solve the systems of equation
x^2=3y+2
x=y-1

Answer by ikleyn(52765)

.

Simply substitute  (y-1) into the first equation instead of x. You will get


    (y-1)^2 = 3y + 2.


Simplify


    y^2 - 2y + 1 = 3y + 2

    y^2 - 5y - 1 = 0.


Next apply the quadratic formula and find two values for y.


Having y, calculate respective values for x.


There are two solution --- two intersection points.

Having my instructions, complete the rest on your own.